# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v1

@@ 5.2 Attack on Polynomial Encoding
1:procedure $\mathsf{PE\_Attack}$($\{\mathsf{L}_{i}(0)\},\mathsf{F}(Y)$) $\triangleright$ $\{\mathsf{L}_{i}(Y)\}_{1\leq i\leq k}$ are input points, $\mathsf{F}(Y)$ is a honest result.
2:     $\textsf{ct}{}_{\sf forged}\leftarrow\textsf{Eval}_{g}(\mathsf{L}_{1}(0),\dots,\mathsf{L}_{k}(0))$triangleright $g$ is any circuit different from $f$.
3:     $\mathsf{L}^{\star}(Y)\leftarrow\textsf{Enc}(u_{1},\dots,u_{N})Y$triangleright Every $u_{i}$ is a unit.
4:     $\mathsf{P}(Y)\leftarrow p({\sf L}^{\star}(Y))$triangleright $p(y)=y^{\phi(t)}$.
5:     $\mathsf{H}(Y)\leftarrow\textsf{ct}{}_{\sf forged}+(\mathsf{F}(Y)-\textsf{ct}{}_{\sf forged})\cdot\mathsf{P}(Y)$
6:     return $\mathsf{H}(Y)$
7:end procedure
Now we have the following theorem.
@@
PE is not secure against the presence of $\mathcal{A}_{c}^{\mathcal{O}_{\sf ReQ}}.$
6 Implementation
We implemented our attacks on the Oblivious Ride Hailing implementation in the Veritas library, which is based on the Lattigo BFV Implementation [MBTPH20]. Specifically, we implemented the attack on REP (Algorithm 3) as well as the attack on $\textsf{REP}{}^{\sf MSK}$ (Algorithm 4). Since the Lattigo library currently does not support BFV bootstrapping, we modified the value of $t$ in the parameters provided by the example code in the Veritas library to ensure that the forgery circuit operates without bootstrapping. However, when bootstrapping is supported, the attack remains feasible regardless of the value of $t$  [^3] For PE, $t$ must be of size roughly equal to the security parameter $\lambda$, since the secret $\alpha$ is chosen from $\mathbb{Z}_{t}^{\times}$. With this parameter setting, the implementation of the attack is not feasible without bootstrapping. Thus we did not implement this attack on PE..
@@ 6.1 Attack Results
The experiments were performed on an Intel(R) Xeon(R) Silver 4114 CPU at 2.20GHz running Linux in a single-threaded environment. The detailed parameters and the corresponding attack results are summarized in Tables 333. The time cost only measured evaluation time. We note that the attack cost is the difference between the evaluation costs of the honest and cheating circuits, not determined by the ratio between them.
Parameters $\log N$ $\log Q$ $t$ $n$
REP Attack Exp1 16 1440 $12^{19}+2^{18}+$ 1 64
$\textsf{REP}{}^{\sf MSK}$ttack Exp2 15 700 $12^{16}+$ 1 64
Exp3 32
@@
Evaluation Time (Sec/op) Honest Circuit Evaluation Cheating Circuit Evaluation Time Difference (Attack Cost)
REP Attack Exp1 2.47s 1417.45s 1414.08s
$\textsf{REP}{}^{\sf MSK}$ttack Exp2 0.31s 5.94s 5.63s
Exp3 0.31s 5.41s 5.10s
Table 2: Circuit Evaluation Cost
@@
REP Attack Exp1 100% 100% 1
$\textsf{REP}{}^{\sf MSK}$ttack Exp2 1.13% 0.80% 1000
Exp3 2.22% 3.60% 500
Table 3: Attack Success Probability
6.1.1 Attack on REP.
In experiment 1, we implement the attack on REP described in 3. We generated keys whose rotation indices are divisors of $n$, assuming a scenario in the adversarial server needs to bootstrap with these rotation keys. To optimize the attack circuit, we modified $\textsf{Eval}_{\sf Interpolate}$ in Algorithm 3 to perform $\textsf{ct}{}\mapsto\textsf{ct}{}-\textsf{Enc}(1,\dots,1)$ and omitted $\textsf{Eval}_{{\sf Normalize}_{r}}$. As a result, we obtain a constant multiple of $J_{S^{c}}$.
This attack circuit can be implemented without bootstrapping under the parameters $(\log N=15,\log Q=700)$. However, due to issues with the large multiplication depth of the circuit, the attack failed with approximately half probability due to the noise budget issue. Nevertheless, this failure is not due to a flaw in the attack itself but rather a limitation arising from the inability to perform bootstrapping. Since this attack assumes a scenario in which the adversarial server possesses $\textsf{rtk}{}_{1}$ for bootstrapping, it does not contradict the result of Corollary 2. Instead, we implemented the attack using parameters $(\log N=16,\log Q=1440)$, which provide a larger noise budget. Due to the large $N$ and the replication number $n$, evaluating the cheating circuit incurs a substantial computational cost. Nevertheless, the attack remains feasible within 30 minutes. Moreover, since Algorithm 3 is inherently parallelizable, leveraging GPU optimization could potentially reduce the runtime to just a few seconds. This assumption is highly realistic, as adversaries in the context of VHE are generally considered to possess extremely high computational power.
6.1.2 Attack on $\textsf{REP}{}^{\sf MSK}.$
In experiments 1 and 2, we implement the attack on $\textsf{REP}{}^{\sf MSK}$ described in Algorithm 4. Since this attack naturally applies to REP, without any modification to the REP scheme we implemented this attack: Specifically, we did not generate the $\textsf{rtk}{}_{1}$. To optimize the attack circuit, we employed the following strategy: Since $\phi(t)=2^{16}$, when constructing the characteristic function in Lemma 3, we can choose $d$ to be a power of 2. In Experiment 2, where $n=64$, we set $d=32\cong\mu=33$ for the characteristic function. According to Theorem 4.2, the corresponding expected attack success probability is $(1-\frac{2^{11}}{2^{16}+1})^{32}-(1-\frac{2^{11}}{2^{16}+1})^{33}\cong 1.13\%$. Also in Experiment 3, where $n=32$, we set $d=16\cong\mu=17$. The corresponding expected attack success probability is $(1-\frac{2^{12}}{2^{16}+1})^{16}-(1-\frac{2^{12}}{2^{16}+1})^{17}\cong 2.22\%$.
@@ 7.1 Homomorphic Cryptanalysis
To provide verifiability, the VHE schemes in [CKP+24, ADDG24] utilized auxiliary secret information as well as the secret key of the HE: a set of indices $S$ in REP and an element $\alpha\in\mathbb{Z}_{t}^{\times}$ in PE. The attack against the schemes described in this paper consists of two steps as follows;
@@ 7.1.2 Step 2: Modify a legitimate ciphertext
This clearly exhibits the structure of the previously discussed cheating circuit.
@@ 8 Conclusion
For the vADDG scheme in [ADDG24], we introduced a shallow pseudo-random characteristic function to overcome the limitation of bootstrapping depth. When a functional bootstrapping with $k$-bit inputs are allowed, the attack probability is $2^{-\beta/(2^{k}\ln 2)}$ with computational cost $O(\gamma)$. For REP in [CKP+24], we constructed a circuit that homomorphically calculates the position vector of the common values. This attack has cost $O(n\log t)$ with probability 1. Our patching solution using multiple secret keys lowers the success probability into $O(n^{-1})$, which provides very weak security, possibly against a covert adversary. For PE, we exploit Euler’s theorem to construct a deterministic forgery attack with computational cost $O(\log t)$.
One meaningful approach to decrease the success probability of attacks was to restrict the circuit depth that is homomorphically evaluated. However, as can be seen in the attack on vADDG, a cheating circuit can still be constructed even at shallow depths. Moreover, an adversary can perform forgery despite risking computation failure due to error robustness. Thus, this approach of limiting depth requires further precise analysis.
@@ Acknowledgement
We would like to thank Damien Stehlé for his guidance and valuable comments on this research.

### v2

@@ 5.2 Attack on Polynomial Encoding
1:Input: $\{\mathsf{L}_{i}(Y)\}_{1\leq i\leq k}$ are input points and $\mathsf{F}(Y)$ is a honest result.
2:Output: ${\sf H}(Y)$, a forged ciphertext polynomial.
3:procedure $\mathsf{PE\_Attack}$($\{\mathsf{L}_{i}(0)\},\mathsf{F}(Y)$)
4:     $\textsf{ct}{}_{\sf forged}\leftarrow\textsf{Eval}_{g}(\mathsf{L}_{1}(0),\dots,\mathsf{L}_{k}(0))$triangleright $g$ is any circuit different from $f$.
5:     $\mathsf{L}^{\star}(Y)\leftarrow\textsf{Enc}(u_{1},\dots,u_{N})Y$triangleright Every $u_{i}$ is a unit.
6:     $\mathsf{P}(Y)\leftarrow p({\sf L}^{\star}(Y))$triangleright $p(y)=y^{\phi(t)}$.
7:     $\mathsf{H}(Y)\leftarrow\textsf{ct}{}_{\sf forged}+(\mathsf{F}(Y)-\textsf{ct}{}_{\sf forged})\cdot\mathsf{P}(Y)$
8:     return $\mathsf{H}(Y)$
9:end procedure
@@
PE is not secure against the presence of $\mathcal{A}_{\sf cov}^{\mathcal{O}_{\sf ReQ}}.$
However, this attack can be detected by counting the number of $\sf ReQ$ calls. The honest output is ${\sf F}(Y)$, and in order to compute $H(Y)$, one must additionally compute ${\sf P}(Y)$ and then multiply it by ${\sf F}(Y)-\textsf{ct}{}_{\sf forged}$. Therefore, compared to the honest circuit, the number of $\sf ReQ$ calls increases by at least $\lceil\log\phi(t)\rceil$. Thus, in this case, the adversary needs to deceive the client about the number of ReQ calls required to evaluate the honest circuit.
6 Attack Implementation on REP and $\textsf{REP}{}^{\sf msk}$
We implemented our attacks in the Veritas library, which is based on the Lattigo BFV Implementation [MBTPH20]. Specifically, we implemented the attack on REP (Algorithm 3) as well as the attack on $\textsf{REP}{}^{\sf msk}$ (Algorithm 4). Since the Lattigo library currently does not support BFV bootstrapping, we modified the value of $t$ in the parameters provided by the example code in the Veritas library to ensure that the forgery circuit operates without bootstrapping. However, when bootstrapping is supported, the attack remains feasible regardless of the value of $t$.  [^4] For PE, $t$ must be of size roughly equal to the security parameter $\lambda$, since the secret $\alpha$ is chosen from $\mathbb{Z}_{t}^{\times}$. With this parameter setting, the implementation of the attack is not feasible without bootstrapping. Thus, we did not implement attack on PE due to the lack of bootstrapping functionality.. Specific parameters are given in table 3.
@@ 6 Attack Implementation on REP and $\textsf{REP}{}^{\sf msk}$
6.1 Attack Description
The attack on REP and $\textsf{REP}{}^{\sf msk}$ aims to implement the following circuit:
$\mathsf{Mult}(\textsf{Eval}_{f}(\textsf{ct}{}),\textsf{Enc}(J_{S}))+c\cdot\textsf{Enc}(J_{S^{c}})$
for some nonzero constant $c$. Since $\textsf{Enc}(J_{S})=(1,\dots,1)-\textsf{Enc}(J_{S^{c}})$, the attack reduces to recovering a scalar multiple of $\textsf{Enc}(J_{S^{c}})$. Therefore, the attack cost corresponds to the amount of resources required to recover a scalar multiple of $\textsf{Enc}(J_{S^{c}})$.
@@ 6.2 Attack Results
The experiments were performed on an Intel(R) Xeon(R) Silver 4114 CPU at 2.20GHz running Linux in a single-threaded environment. The attack results are summarized in Tables 33. The time cost only measured evaluation time for $\textsf{Enc}(J_{S^{c}})$.
6.2.1 Attack on REP.
In experiment 1 to 4, we implement the attack on REP described in Algorithm 3 with various $k=1,2,4,33$. We generated keys whose rotation indices are divisors of $n$, assuming a scenario in the adversarial server needs to bootstrap with these rotation keys. We optimized the attack circuit described in Algorithm 3 as follows. Instead of evaluating $\textsf{Eval}_{\sf Interpolate}$, which maps $\textsf{Enc}(1,\dots,1)$ to $\textsf{Enc}(0,\dots,0)$ and $\textsf{Enc}\left(n/2,\dots,n/2\right)$ to $\textsf{Enc}(1,\dots,1)$, we directly performed the subtraction $\textsf{ct}{}\mapsto\textsf{ct}{}-\textsf{Enc}(1,\dots,1)$. This operation maps $\textsf{Enc}(1,\dots,1)$ to $\textsf{Enc}(0,\dots,0)$ and $\textsf{Enc}\left(n/2,\dots,n/2\right)$ to $\textsf{Enc}\left(n/2-1,\dots,n/2-1\right)$. As a result, we obtain a ciphertext that corresponds to a constant multiple of $J_{S^{c}}$.
According to Theorem 4.2, the expected attack success probabilities given by $1-\binom{32}{k}/\binom{64}{k}$ are $50\%$, $75.40\%$, $94.34\%$, and $100\%$ for $k=1$, $2$, $4$, and $33$, respectively. This attack circuit can be implemented without bootstrapping under the parameters $(\log N=15,\log PQ=880,\log Q=759,t=2^{16}+1)$. As shown in Table 3, the attack time increases linearly with $k$. We note that Algorithm 3 is inherently parallelizable, as it performs ${\sf CVS}(\textsf{ct}{},i)$ and ${\sf is\_in\_set}(\textsf{ct}{},i)$ independently for each $i=1,\dots,k$. Thus, by parallelizing the execution of ${\sf CVS}(\textsf{ct}{},i)$, the overall attack time can be reduced to under one minute even with $k=33$. In addition, leveraging GPU-based optimizations could potentially reduce the runtime to just a few seconds. This assumption is realistic, given that adversaries in the context of HE are typically assumed to have access to significant computational resources.
6.2.2 Attack on $\textsf{REP}{}^{\sf msk}.$
We implement the attack on $\textsf{REP}{}^{\sf msk}$ described in Algorithm 4. Since this attack naturally applies to REP, without any modification to the REP scheme we implemented this attack: Specifically, we did not generate the $\textsf{rtk}{}_{1}$.
Since $\phi(t)=2^{16}$, when constructing the characteristic function in Lemma 2, we can choose $d$ to be a power of 2. In Experiment 5, where $n=64$, we set $d=32\cong\mu=33$ for the characteristic function. As a result, the output becomes a scalar multiple of $\textsf{Enc}(J_{S^{c}})$ with non-negligible probability. According to Theorem 4.3, the corresponding expected attack success probability is $\left(1-\frac{2^{11}}{2^{16}+1}\right)^{32}-\left(1-\frac{2^{11}}{2^{16}+1}\right)^{33}\cong 1.13\%$. Also in Experiment 6, where $n=32$, we set $d=16\cong\mu=17$. The corresponding expected attack success probability is $\left(1-\frac{2^{12}}{2^{16}+1}\right)^{16}-\left(1-\frac{2^{12}}{2^{16}+1}\right)^{17}\cong 2.22\%$.
Parameters $\log N$ $\log PQ$ $\log Q$ $t$ $|S|=|\{1,\dots,n\}|$ $k$
REP Attack Exp1 15 880 759 $12^{16}+$ 1 64 1
Exp2 2
Exp3 4
Exp4 33
$\textsf{REP}{}^{\sf msk}$ttack Exp5 15 880 759 $12^{16}+$ 1 64 -
Exp6 32 -
@@
Evaluation Time (Sec/op) Attack Time (Evaluating $\textsf{Enc}(J_{S^{c}})$) Remark
REP Attack Exp1 8.37s $\textsf{rtk}{}_{1}$ generated.
Exp2 15.91s
Exp3 29.67s
Exp4 246.54s
$\textsf{REP}{}^{\sf msk}$ttack Exp5 6.05s -
Exp6 6.36s -
Table 2: Attack Circuit Evaluation Cost
@@
REP Attack Exp1 50.00% 51.8% 1000
Exp2 75.40% 75.5% 1000
Exp3 94.34% 93.4% 1000
Exp4 100% 100% 1
$\textsf{REP}{}^{\sf msk}$ttack Exp5 1.13% 0.8% 1000
Exp6 2.22% 1.8% 1000
Table 3: Attack Success Probability (1000 iterations unless otherwise noted)
Due to the lack of BFV bootstrapping functionality in the Lattigo library, we implemented the attacks on REP and $\textsf{REP}{}^{\sf msk}$ only for specific parameter settings, and did not implement the attack on PE. However, we note that in a bootstrappable setting, not only is an attack on PE feasible, but the above attacks can also be carried out independently of the parameter settings.
@@ 7.1 Homomorphic Cryptanalysis
To provide verifiability, the VHE schemes in [CKP+24, ADDG24] utilized auxiliary secret information as well as the secret key of the HE: a set of indices $S$ in REP and an element $\alpha\in\mathbb{Z}_{t}^{\times}$ in PE. Attacks against the schemes described in this paper consist of two steps as follows;
@@ 7.1.2 Step 2: Modify a legitimate ciphertext
This exhibits the structure of the previously discussed cheating circuit.
@@ 8 Conclusion
For the vADDG scheme in [ADDG24], we introduced a shallow pseudo-random characteristic function to overcome the limitation of bootstrapping depth. The attack probability is $\left(1-\left(1/2\right)^{\alpha}\right)\cdot\left(1/2\right)^{\beta}$ with computational cost $O(\gamma)$. For REP in [CKP+24], we constructed a circuit that homomorphically calculates the position vector of the common values. This attack has cost $O(k\log t)$ with probability 1 when $k=n/2+1$. Our patching solution using multiple secret keys lowers the success probability into $O(n^{-1})$, which provides very weak security, possibly against a covert adversary. For PE, we exploit Euler’s theorem to construct a deterministic forgery attack with computational cost $O(\log t)$.
One meaningful approach to decrease the success probability of attacks was to restrict the circuit depth that is homomorphically evaluated. However, as seen in the attack on vADDG, a cheating circuit can still be constructed even at shallow depths. Moreover, an adversary can perform forgery despite risking computation failure due to error robustness. Thus, this approach of limiting depth requires further precise analysis.
@@ Acknowledgement
We would like to thank Damien Stehlé and Changmin Lee for his guidance and valuable comments on this research.